A particle starts from the origin at t = 0 with an initial velocity having an x component of 20 m/s and a y component of -15 m/s. The particle moves in the xy plane with an x component of acceleration only, given by ax = 4.0 m/s2. (a) Determine the components of the velocity vector at any time and the total velocity vector at any time. (b) Calculate the velocity and speed of the particle at t = 5.0 s. (c) Determine the x and y coordinates of the particle at any time t and the position vector at this time.
1
Expert's answer
2020-02-10T09:30:41-0500

Solution. According to the conditions of the problem for t=0s


"x_0=0m"

"y_0=0m""v_{0x}=20 \frac{m}{s}"

"v_{0y}=-15 \frac{m}{s}""a_x=4 \frac{m}{s^2}"

"a_y=0 \frac{m}{s^2}"

a) The particle moves with constant acceleration along the 0x axis; therefore, the component vx can be written as


"v_x(t)=20+4t \frac{m}{s}"

The particle moves at a constant speed along the axis 0y; therefore, the component wu can be written as


"v_y(t)=-15 \frac{m}{s}"

Therefore, the total velocity vector is equal to

"v(t)=(20+4t)i-15j"

b)  Calculate the velocity at t = 5.0 s.


"v(5)=(20+20)i-15j=40i-15j"

Calculate the speed at t = 5.0 s.


"v=\sqrt{20^2+(-15)^2}=\sqrt{625}=25 \frac{m}{s}"

c) Integrating the velocity equations of the components by the time and using the values of the initial coordinates, we obtain


"x(t)=20t+2t^2""y(t)=-15t"

As result position vector at this time


"r(t)=(20t+2t^2)i-15tj"

Answer. a)


"v_x(t)=20+4t \frac{m}{s}""v_y(t)=-15 \frac{m}{s}""v(t)=(20+4t)i-15j"

b) velocity


"v(5)=40i-15j"

speed


"v=25 \frac{m}{s}"

c)


"x(t)=20t+2t^2"

"y(t)=-15t"

"r(t)=(20t+2t^2)i-15tj"